OpenStudy (anonymous):

Please Help! Given the quadratic equation: x2 - 8x + 11 = 0 Write the quadratic function in vertex form.

OpenStudy (anonymous):

To get a quadratic in vertex form you need to "complete the square". This entails rearranging it into the form of a(x-b)^2+c So we put the x^2 and -8x in a bracket \[(x^2-8x)+11=0\] Now we are going to move the x's square outside the bracket, but first we need to half the -8 and take it's x off so if we were to expand the bracket later it would form a -8x again. \[(x-4)^{2}+11=0\] If we were to expand that bracket now sure we would get our -8x, but we would also get a random 16 popping out that wasn't there before, so to counter act this we tag -16 onto the end\[(x-4)^{2}+11-16=0\]Which simplifies to\[(x-4)^{2}-5=0\] giving us our answer :)

OpenStudy (anonymous):

y = x² + 8x + 11 line of symmetry is x = -8/2 = -4 y = 4² + 8(-4) + 11 = 16 – 32 + 11 = -5 so vertex is (-4, -5) since x axis is 5 up from vertex, and for quads, up is over squared, zeros at -4 ± √5 x² + 8x + 11 = 0 x² + 8x + ... = -11 ∙ ∙ ∙ ∙ ∙ ∙ add (-8/2)² x² + 8x + 16 = -11 + 16 (x + 4)² = 5 x + 4 = ±√5 x = -4 ± √5